# Talent Hoarding in Organizations

### v2

@@ 5.3 Discussion and Interpretation
The preceding results provide empirical evidence that the worker application responses to manager rotations are primarily driven by managerial talent hoarding. Although it is not possible to exhaustively rule out all potential alternative mechanisms, the observed patterns impose strong restrictions on the manner in which such mechanisms would have to affect worker behavior, which provides evidence against a quantitatively meaningful role for many plausible alternative mechanisms. Nonetheless, it remains possible that the observed responses to rotations may also capture factors that are unrelated to talent hoarding. Yet, to the extent that talent hoarding is a key driver of responses to manager rotations, analyzing impacts of rotations is informative about the impacts of hoarding on workers’ careers. This interpretation is supported by the analysis focusing on managers in Section 4. The following section analyzes impacts on workers’ subsequent career outcomes.
@@ 6.1 IV Approach
I estimate $\beta_{IV}$ by two-stage least squares, which can be interpreted as the local average treatment effect (LATE), defined as the effect of applications on hiring outcomes for workers induced to apply by manager rotations. Because workers can only get hired if they apply and there are no always takers, LATE equals the treatment effect on the treated (TOT). To interpret $\beta_{IV}$ as LATE, four assumptions are required to hold: relevance, independence, exclusion, and monotonicity (Angrist and Imbens, 1995, Angrist et al., 1996). I now provide evidence in support of these assumptions. Relevance.— Rotations have large and significant effects on worker applications (Section 5). Independence.— Manager and worker characteristics are balanced and neither worker applications nor other key team-level outcomes, such as absenteeism or bonus pay, are characterized by pre-trends, suggesting that rotations are as good as randomly assigned (Section 5). Exclusion.—The exclusion restriction requires that manager rotation does not affect hiring outcomes other than through workers’ decisions to apply. This assumption would be violated if departing managers intercede in workers’ hiring outcomes, for instance by taking workers with them to their new team or replacing themselves with workers. However, this occurs in less than 3% of applications, suggesting that departing managers generally do not make subsequent hiring decisions for their subordinates. Alternatively, departing managers may influence hiring outcomes by reaching out to hiring managers. I test for this possibility by comparing applications where the posted job is located in the same area as workers’ current job, making it more likely that the departing manager interacts with the hiring manager. However, I do not find that relative hiring rates are higher for job openings to which managers have closer ties (Panel A of Appendix Table B7). Another violation could occur if manager rotation increases worker qualifications, making them more likely to get hired. If this channel is quantitatively important, one would expect to see larger hiring effects for workers under managers with higher quality or who have been exposed to the manager for longer. I measure manager quality using past means of three outcomes: team turnover, team absenteeism, and managers’ performance rating. Panel B of Appendix Table B7 documents that the marginal hiring probability is similar for high-quality vs low-quality managers. Panel C of Appendix Table B7 shows that marginal hiring probabilities are comparable across exposure length, even for workers who have been with the manager for only one quarter, suggesting that such a channel is unlikely to violate the exclusion restriction. Monotonicity.— My results do not appear to be biased by the existence of a large defier population. In my setting, defiers are individuals who would have applied in the absence of a manager rotation, but whose unobserved propensity to apply is reduced by the instrument. Following Arnold et al. (2018) and Bhuller et al. (2020), I show that the first-stage relationship between applications and manager rotations remains positive for 16 key subgroups of workers (Panel A of Appendix Table B8). To further test for the presence of defiers, I use two proxies for workers’ unobserved application propensity: workers’ own past application behavior and the leave-out team mean of past application rates. Panel B of Appendix Table B8 shows that even workers, for whom the proxies predict that they are more likely to be defiers because of their higher application propensity in the past, do not experience lower or negative application effects when managers rotate.
@@ 6.2 Impacts on the Allocation of Talent
A complier analysis based on Abadie (2003) corroborates that talent hoarding reduces applicant quality. [^26] Under standard IV assumptions, complier characteristics can be estimated as $E[X_{it}|\text{Compliers}]$ for some characteristic $X_{it}$. I calculate average complier characteristics and standard errors by performing 2SLS using the first-stage Equation 3 and a reduced-form equation replacing the outcome variable in Equation 4 with $X_{it}A_{it}$, where $X_{it}$ corresponds to a characteristic of individual $i$ and $A_{it}$ is a binary indicator for $i$ applying in quarter $t$. I compute characteristics for always takers, who apply even in the presence of talent hoarding, by estimating an OLS regression of $X_{it}A_{it}(1-Z_{it})$ on $A_{it}(1-Z_{it})$, which allows me to estimate $E[X_{it}|\text{Always takers}]$. Panel A of Table 6 compares average characteristics across the entire employee population (Column 1), always takers who apply even in absence of manager rotations (Column 2), and marginal applicants who only apply if a manager rotates (Column 3). I find that marginal applicants are positively selected. While 48.9% of always takers hold a graduate degree, this is true for 59.4% of marginal applicants. Similarly, 61.8% of always takers received high performance ratings prior to applying, compared to 70.4% of marginal applicants. These results are consistent with the predictions of the conceptual framework, which indicate that talent hoarding reduces the share of high-quality workers in the applicant pool, potentially limiting the ability of the firm to fill positions with suitable candidates. The fact that marginal applicants are of higher quality and that the firm would have liked to hire a high share of them suggests that these applicants would have been good fits for the new positions. To provide suggestive evidence that these applicants would have likely been productive in their new positions, I use data on employees’ performance ratings, which are designed to provide task-specific feedback on whether a worker has accomplished her tasks in the past evaluation cycle. Since most workers in a team perform very similar tasks, performance ratings are particularly well-suited for drawing comparisons across workers within teams. I assess the possibility of forgone performance by estimating a 2SLS regression in which the outcome is defined as a worker landing a position and performing better than the average performance of other workers in the new team one year later. Column 3 of Table 4, Panel A presents the results of this analysis and shows that 17.7% of marginal applicants who apply to and land a new position also outperform their teammates.
@@
In addition to creating efficiency costs, talent hoarding may impact between-group inequality if some workers are more affected than others. The survey evidence in Section 2.1 indicates that talent hoarding occurs through direct interpersonal interactions, suggesting that impacts may be larger for workers who depend more heavily on managerial support or are more sensitive to confrontation with their manager. In line with previous work (Bertrand, 2011), employees’ survey responses highlight the possibility that talent hoarding may affect women more than men (Figure 6). Women are 26% more likely to mention the importance of manager support for their career development ($p$=0.000) and are 22% more likely to rank a good relationship with their supervisor as the most important feature of their job ($p$=0.001). These differences suggest that women rely more on managers’ career guidance and place more value on preserving a good relationship with their manager. I conduct three exercises to assess the extent of gender differences. First, I conduct the complier analysis separately by gender in order to analyze whether talent hoarding affects women at a higher part of the quality distribution. Panels B and C of Table 6 compare average characteristics across the entire employee population (Columns 4 and Column 7), always takers who apply even in absence of manager rotations (Columns 5 and 8), and marginal applicants who only apply if a manager rotates (Columns 6 and Column 9). For both men and women, marginal applicants are positively selected. However, the extent of this positive selection is more pronounced for women. While 61.2% of marginal female applicants hold a graduate degree, this is only true for 36.3% of always takers. For men, 59.4% of marginal applicants and 52.5% of always takers hold a graduate degree. Similarly, 77.7% of women at the margin received high performance ratings relative to 61.5% of always takers. Among men, 69.2% of marginal applicants versus 62.0% of always takers received high ratings. These patterns suggest that higher-quality women are affected by talent hoarding. Second, I compare mobility effects by gender by estimating the hiring probabilities for men and women at the margin. Columns 3 and 4 of Table 5 report the 2SLS estimates of Equation 5, which are separately estimated by gender and capture the probability of landing a higher-level position for marginal applicants. I find that both men and women experience positive and statistically significant marginal hiring probabilities of 41.2% and 62.3%, respectively. These hiring probabilities are much larger than those of always takers in my sample, which are 26.9% for men and 28.6% for women. Although the gender difference among marginal applicants is not statistically significant, the extent of positive selection within women is striking (i.e., marginal vs. average applicants), and suggests that talent hoarding may be more consequential for women’s career outcomes. A third test for disparate impacts by gender compares the average potential outcomes with respect to compliers’ earnings in the treated state (i.e., when marginal applicants apply because talent hoarding ceases) to those in the untreated state (i.e., when marginal applicants do not apply due to talent hoarding). This analysis leverages the potential outcomes framework that follows from the interpretation of $\beta_{IV}$ as the LATE for marginal applicants. The advantage of this approach is that it allows a comparison of the same set of individuals across two different potential outcomes, avoiding potential composition bias. Figure 7 presents estimates of log annual earnings in quarter $t+4$ for marginal applicants in both potential outcomes states by gender. Outcomes are reported in terms of percentiles at the firm. Both men and women experience higher earnings if they choose to apply; however, the larger gains realized by women lead to a reduction in the pay gap of 86%. These findings suggest that talent hoarding may exacerbate gender inequality in the firm, raising the possibility that talent hoarding has non-negligible consequences with respect to both efficiency and equity in the internal labor market.
@@ 7 Conclusion
This paper provides the first empirical evidence that talent hoarding is an important source of frictions in organizations. Using a unique combination of survey and administrative data from a large manufacturing firm, I show that managers face strong incentives to hoard talented workers in their team and act accordingly. My findings indicate that managerial talent hoarding is commonplace (explicitly reported by 75% of managers) and has non-negligible effects on the internal allocation of talent within the firm. While my results provide the first detailed insights on talent hoarding, additional evidence suggests that such talent hoarding behavior is endemic. In a survey of the top publicly listed companies in Germany, 83% cite talent hoarding as a crucial friction in their organization (hkp, 2021). Firm surveys in other countries, such as the United States, document that talent hoarding is not limited to German organizations (i4cp, 2016, KornFerry, 2015). Because talent hoarding arises due to misaligned incentives, a natural solution would be to more closely align the incentives of managers with those of the firm. In theory, firms could reward managers based on their subordinates’ future performance, irrespective of whether subordinates remain on the team. However, HR leaders typically state that realigning incentives through financial compensation is infeasible (hkp, 2021). Similarly, when asked under which conditions managers would be more likely to support workers’ career development, only a minority of managers cite financial incentives. Instead, 68% of managers highlight the ease of replacing lost talent as key condition. In addition, policies that increase application rates — such as implementing regular application schedules and having other organizational agents, such as the HR department, directly invite workers to apply for positions — could reduce the scope for managers to engage in talent hoarding. The survey I conduct shows that for such policies to be effective, the firm must be able to deter managers from retaliating against workers, for instance by assuring full confidentiality for applicants.
@@
Appendix A Appendix Figures - For Online Publication
Figure A1: Manager-Reported Incentives and Actions Regarding Talent Development
Panel A. Incentives for Managers to Support Workers’ Careers
Panel B. Effective Ways for Managers to Impact Workers’ Careers
Notes: This figure uses responses from the manager survey to depict manager-reported incentives and actions with respect to talent development. Panel A shows the share of managers who indicated that a given condition would make managers more likely to support workers’ career development. Panel B shows the share of managers who chose a given action as an effective way through which managers can impact their workers’ career development. Managers were allowed to choose multiple answers. See Appendix Section D.3 for more information on the manager survey and the exact question wording. 95%-level confidence intervals are displayed. Controls: Female, age, German citizenship, firm tenure, function. N=3,xxx.
Panel B. Effective Ways for Managers to Impact Workers’ Careers
@@ Appendix A Appendix Figures - For Online Publication
Figure A3: Effect of Manager Rotations by Workers’ Exposure Length to Manager
@@
Figure A5: Placebo Tests for Effects of Manager Rotations
Panel A. Results by Outcome of Manager Application
Panel B. Results by Type of Rotating Teammate
Panel B. Results by Type of Rotating Teammate
Appendix B Appendix Tables - For Online Publication
@@
Form of talent hoarding
Quotations by workers
Informal underrating
“[My] former boss made me look bad at the potential new superiors behind my back in order to keep me on the team.”
Formal underrating
“Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.”
@@
Suppressed development
“[Supervisors] keep employees in their positions by preventing further development, rejecting training courses, and increasing workloads to prevent capacity for new tasks and development measures.”
@@
Soft pressure
“I decided not to apply internally … because the message communicated to me in the employee dialogue was that I can’t leave the team within three years of joining.”
@@
Threats
“… my supervisor communicated very openly in a workshop that if one of his employees applied and was not hired, his career in his current department would also be at an end.”
Retaliation
“The position should fit exactly so that negative effects of applying regarding the current manager are worthwhile.”
@@
Statement
Quotations by managers
Misaligned incentives
“Regarding the development of your subordinates, there often is a conflict of interests for managers, since the employee then usually leaves the team and the position will not be approved to be refilled again.”
@@
Risk of losing talent
“I observe that managers are not interested in letting good workers leave their current position or develop them, otherwise they would have to fill the current position again.”
@@
Appendix C Theoretical Appendix - For Online Publication

### v3

@@ 5.3 Discussion and Interpretation
The preceding results provide empirical evidence that the worker application responses to manager rotations are to a large part driven by managerial talent hoarding. Appendix Section E.3.5 complements this interpretation by providing evidence against several plausible alternative mechanisms—including bad-news shocks, milestone completions, loyalty, match quality, and role-model effects—as the primary drivers of the observed effects. Although it is not possible to exhaustively rule out all potential alternative mechanisms, the observed patterns impose strong restrictions on the manner in which such mechanisms would have to affect worker behavior, which provides evidence against a quantitatively meaningful role for many plausible alternative mechanisms. Nonetheless, it remains possible that the observed responses to rotations may also capture factors that are unrelated to talent hoarding. Yet, to the extent that talent hoarding is a key driver of responses to manager rotations, analyzing impacts of rotations is informative about the impacts of hoarding on workers’ careers. This interpretation is supported by the analysis focusing on managers in Section 4. The following section analyzes impacts on workers’ subsequent career outcomes.
@@ 6.1 IV Approach
I estimate $\beta_{IV}$ by two-stage least squares, which can be interpreted as the local average treatment effect (LATE), defined as the effect of applications on hiring outcomes for workers induced to apply by manager rotations. Because workers can only get hired if they apply and there are no always takers, LATE equals the treatment effect on the treated (TOT). To interpret $\beta_{IV}$ as LATE, four assumptions are required to hold: relevance, independence, exclusion, and monotonicity (Angrist and Imbens, 1995, Angrist et al., 1996). I now provide evidence in support of these assumptions.
Relevance.— Rotations have large and significant effects on worker applications (Section 5).
Independence.— Manager and worker characteristics are balanced and neither worker applications nor other key team-level outcomes, such as absenteeism or bonus pay, are characterized by pre-trends, suggesting that rotations are as good as randomly assigned (Section 5).
Exclusion.—The exclusion restriction requires that manager rotation does not affect hiring outcomes other than through workers’ decisions to apply. This assumption would be violated if departing managers intercede in workers’ hiring outcomes, for instance by taking workers with them to their new team or replacing themselves with workers. However, this occurs in less than 3% of applications, suggesting that departing managers generally do not make subsequent hiring decisions for their subordinates. Alternatively, departing managers may influence hiring outcomes by reaching out to hiring managers. I test for this possibility by comparing applications where the posted job is located in the same area as workers’ current job, making it more likely that the departing manager interacts with the hiring manager. However, I do not find that relative hiring rates are higher for job openings to which managers have closer ties (Panel A of Appendix Table B8).
Another violation could occur if manager rotation increases worker qualifications, making them more likely to get hired. If this channel is quantitatively important, one would expect to see larger hiring effects for workers under managers with higher quality or who have been exposed to the manager for longer. I measure manager quality using past means of three outcomes: team turnover, team absenteeism, and managers’ performance rating. Panel B of Appendix Table B8 documents that the marginal hiring probability is similar for high-quality vs low-quality managers. Panel C of Appendix Table B8 shows that marginal hiring probabilities are comparable across exposure length, even for workers who have been with the manager for only one quarter, suggesting that such a channel is unlikely to violate the exclusion restriction.
Monotonicity.— My results do not appear to be biased by the existence of a large defier population. In my setting, defiers are individuals who would have applied in the absence of a manager rotation, but whose unobserved propensity to apply is reduced by the instrument. Following Arnold et al. (2018) and Bhuller et al. (2020), I show that the first-stage relationship between applications and manager rotations remains positive for 16 key subgroups of workers (Panel A of Appendix Table B9). To further test for the presence of defiers, I use two proxies for workers’ unobserved application propensity: workers’ own past application behavior and the leave-out team mean of past application rates. Panel B of Appendix Table B9 shows that even workers, for whom the proxies predict that they are more likely to be defiers because of their higher application propensity in the past, do not experience lower or negative application effects when managers rotate.
@@ 6.2 Impacts on the Allocation of Talent
A complier analysis based on Abadie (2003) corroborates that talent hoarding reduces applicant quality. [^27] Under standard IV assumptions, complier characteristics can be estimated as $E[X_{it}|\text{Compliers}]$ for some characteristic $X_{it}$. I calculate average complier characteristics and standard errors by performing 2SLS using the first-stage Equation 3 and a reduced-form equation replacing the outcome variable in Equation 4 with $X_{it}A_{it}$, where $X_{it}$ corresponds to a characteristic of individual $i$ and $A_{it}$ is a binary indicator for $i$ applying in quarter $t$. I compute characteristics for always takers, who apply even in the presence of talent hoarding, by estimating an OLS regression of $X_{it}A_{it}(1-Z_{it})$ on $A_{it}(1-Z_{it})$, which allows me to estimate $E[X_{it}|\text{Always takers}]$. Panel A of Table 7 compares average characteristics across the entire employee population (Column 1), always takers who apply even in absence of manager rotations (Column 2), and marginal applicants who only apply if a manager rotates (Column 3). I find that marginal applicants are positively selected. While 48.9% of always takers hold a graduate degree, this is true for 59.4% of marginal applicants. Similarly, 61.8% of always takers received high performance ratings prior to applying, compared to 70.4% of marginal applicants. These results are consistent with the predictions of the conceptual framework, which indicate that talent hoarding reduces the share of high-quality workers in the applicant pool, potentially limiting the ability of the firm to fill positions with suitable candidates.
The fact that marginal applicants are of higher quality and that the firm would have liked to hire a high share of them suggests that these applicants would have been good fits for the new positions. To provide suggestive evidence that these applicants would have likely been productive in their new positions, I use data on employees’ performance ratings, which are designed to provide task-specific feedback on whether a worker has accomplished her tasks in the past evaluation cycle. Since most workers in a team perform very similar tasks, performance ratings are particularly well-suited for drawing comparisons across workers within teams. I assess the possibility of forgone performance by estimating a 2SLS regression in which the outcome is defined as a worker landing a position and performing better than the average performance of other workers in the new team one year later. Column 3 of Table 5, Panel A presents the results of this analysis and shows that 17.7% of marginal applicants who apply to and land a new position also outperform their teammates.
@@
In addition to creating efficiency costs, talent hoarding may impact between-group inequality if some workers are more affected than others. The survey evidence in Section 2.1 indicates that talent hoarding occurs through direct interpersonal interactions, suggesting that impacts may be larger for workers who depend more heavily on managerial support or are more sensitive to confrontation with their manager. In line with previous work (Bertrand, 2011), employees’ survey responses highlight the possibility that talent hoarding may affect women more than men (Figure 6). Women are 26% more likely to mention the importance of manager support for their career development ($p$ = 0.000) and are 22% more likely to rank a good relationship with their supervisor as the most important feature of their job ($p$ = 0.001). These differences suggest that women rely more on managers’ career guidance and place more value on preserving a good relationship with their manager.
I conduct three exercises to assess the extent of gender differences. First, I conduct the complier analysis separately by gender in order to analyze whether talent hoarding affects women at a higher part of the quality distribution. Panels B and C of Table 7 compare average characteristics across the entire employee population (Columns 4 and Column 7), always takers who apply even in absence of manager rotations (Columns 5 and 8), and marginal applicants who only apply if a manager rotates (Columns 6 and Column 9). For both men and women, marginal applicants are positively selected. However, the extent of this positive selection is more pronounced for women. While 61.2% of marginal female applicants hold a graduate degree, this is only true for 36.3% of always takers. For men, 59.4% of marginal applicants and 52.5% of always takers hold a graduate degree. Similarly, 77.7% of women at the margin received high performance ratings relative to 61.5% of always takers. Among men, 69.2% of marginal applicants versus 62.0% of always takers received high ratings. These patterns suggest that higher-quality women are affected by talent hoarding.
Second, I compare mobility effects by gender by estimating the hiring probabilities for men and women at the margin. Columns 3 and 4 of Table 6 report the 2SLS estimates of Equation 5, which are separately estimated by gender and capture the probability of landing a higher-level position for marginal applicants. I find that both men and women experience positive and statistically significant marginal hiring probabilities of 41.2% and 62.3%, respectively. These hiring probabilities are much larger than those of always takers in my sample, which are 26.9% for men and 28.6% for women. Although the gender difference among marginal applicants is not statistically significant, the extent of positive selection within women is striking (i.e., marginal vs. average applicants), and suggests that talent hoarding may be more consequential for women’s career outcomes.
A third test for disparate impacts by gender compares the average potential outcomes with respect to compliers’ earnings in the treated state (i.e., when marginal applicants apply because talent hoarding ceases) to those in the untreated state (i.e., when marginal applicants do not apply due to talent hoarding). This analysis leverages the potential outcomes framework that follows from the interpretation of $\beta_{IV}$ as the LATE for marginal applicants. The advantage of this approach is that it allows a comparison of the same set of individuals across two different potential outcomes, avoiding potential composition bias. Figure 7 presents estimates of log annual earnings in quarter $t+4$ for marginal applicants in both potential outcomes states by gender. Outcomes are reported in terms of percentiles at the firm. Both men and women experience higher earnings if they choose to apply; however, the larger gains realized by women lead to a reduction in the pay gap of 86%. These findings suggest that talent hoarding may exacerbate gender inequality in the firm, raising the possibility that talent hoarding has non-negligible consequences with respect to both efficiency and equity in the internal labor market.
@@ 7 Conclusion
This paper provides the first empirical evidence that talent hoarding is an important source of frictions in organizations. Using a unique combination of survey and administrative data from a large manufacturing firm, I show that managers face strong incentives to hoard talented workers in their team and act accordingly. My findings indicate that managerial talent hoarding is commonplace (explicitly reported by 75% of managers) and has non-negligible effects on the internal allocation of talent within the firm. While my results provide the first detailed insights on talent hoarding, additional evidence suggests that such talent hoarding behavior is endemic. In a survey of the top publicly listed companies in Germany, 83% cite talent hoarding as a crucial friction in their organization (hkp, 2021). Firm surveys in other countries, such as the United States, document that talent hoarding is not limited to German organizations (i4cp, 2016, KornFerry, 2015).
Because talent hoarding arises due to misaligned incentives, a natural solution would be to more closely align the incentives of managers with those of the firm. In theory, firms could reward managers based on their subordinates’ future performance, irrespective of whether subordinates remain on the team. However, HR leaders typically state that realigning incentives through financial compensation is infeasible (hkp, 2021). Similarly, when asked under which conditions managers would be more likely to support workers’ career development, only a minority of managers cite financial incentives. Instead, 68% of managers highlight the ease of replacing lost talent as key condition. In addition, policies that increase application rates—such as implementing regular application schedules and having other organizational agents, such as the HR department, directly invite workers to apply for positions—could reduce the scope for managers to engage in talent hoarding. The survey I conduct shows that for such policies to be effective, the firm must be able to deter managers from retaliating against workers, for instance by assuring full confidentiality for applicants.
@@
Appendix A Appendix Figures
@@
Appendix B Appendix Tables
@@
Form of talent hoarding Quotations by workers
Informal underrating “[My] former boss made me look bad at the potential new superiors behind my back in order to keep me on the team.”
Formal underrating “Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.”
@@
Suppressed development “[Supervisors] keep employees in their positions by preventing further development, rejecting training courses, and increasing workloads to prevent capacity for new tasks and development measures.”
@@
Soft pressure “I decided not to apply internally … because the message communicated to me in the employee dialogue was that I can’t leave the team within three years of joining.”
@@
Threats “… my supervisor communicated very openly in a workshop that if one of his employees applied and was not hired, his career in his current department would also be at an end.”
Retaliation “The position should fit exactly so that negative effects of applying regarding the current manager are worthwhile.”
@@
Statement Quotations by managers
Misaligned incentives “Regarding the development of your subordinates, there often is a conflict of interests for managers, since the employee then usually leaves the team and the position will not be approved to be refilled again.”
@@
Risk of losing talent “I observe that managers are not interested in letting good workers leave their current position or develop them, otherwise they would have to fill the current position again.”
@@
Table B4: Heterogeneity in Manager Hoarding by Incentives Using an Alternative Measure Based on Suppression of Training Opportunities
Outcome variable Top tercile training measure
(1) (2) (3)
Financial incentive (share) 0.2007
Team size -0.0143
Visibility 0.0298
Outcome Mean 0.3001 0.3001 0.3001
Predicted difference p90-p10 0.1380 -0.1403 0.0298
Observations 6,xxx 6,xxx 6,xxx
Notes: This table uses an alternative measure of managers’ hoarding actions based on managers’ withholding of access to high-visibility training opportunities to examine whether managerial talent hoarding is more prevalent among managers with stronger incentives to hoard talent. I use three distinct measures of hoarding incentives. Column 1 captures financial incentives using the share of total compensation that is performance-related. Column 2 uses the number of team members overseen by the manager. Column 3 includes a binary indicator for low talent visibility, defined as working in a functional area where individual achievements are less visible. Each coefficient is estimated separately based on a linear model for managers’ hoarding propensity using an OLS regression of the indicator that managers are hoarding-prone, defined as being in the top tercile of the distribution of mean deviations between predicted and actual training assignments, on one incentive measure at a time. The predicted difference p90-p10 refers to the predicted difference in the outcome between the 90th and 10th percentiles of the continuous regressors (between values 1 and 0 for binary indicators), holding all other covariates at their mean values. See Appendix Section E.2.3 for details on the construction of this alternative hoarding measure. See Appendix Section E.2.1 for definitions of the incentive measures and Appendix Table E3 for their distribution in the sample. Appendix Tables E6 and E7 present complementary robustness results using alternative specifications of the regressors and outcome variable. See Table 2 for my main results using managers’ suppression of potential ratings as baseline measure of hoarding behavior.Controls: Female, age, German citizenship, firm tenure, educational qualification. Robust standard errors in parentheses.
@@
Appendix C Theoretical Appendix